OpenStudy (anonymous):

(2-9c)(-8) A.-16+72c B. 16+72c C. -16-72c D. 16-72c

OpenStudy (mathstudent55):

This is a problem using the distributive property. Do you know how to use the distributive property?

OpenStudy (anonymous):

somewhat

OpenStudy (anonymous):

72c-16 is the answer

OpenStudy (mathstudent55):

Ok, let's go over it. Usually, you see the distributive property used in a problem in which the number is to the left of the parentheses. Let's do that: \((2-9c)(-8)\) \(= (-8)(2-9c)\) Here's an example of the distributive property to remind you of how it is used: \(2(3 + 4) = 2 \times 3 + 2 \times 4\)

OpenStudy (mathstudent55):

Now we do the same using the expression in your problem: \(= (-8)(2-9c)\) \(= (-8)(2) - (-8)(9c) \) \(= -16 - (-72c)\) \(= -16 + 72c\)